# Black-Box On-Policy Distillation of Large Language Models

### v2

@@ A.1 Implement GAD with GRPO
We implement policy optimization of the student with GRPO [grpo]. We use $q_{G}$ to denote the output distribution of student $G$. For each input prompt $x$, we sample a group of $N$ student responses $\{y_{s}^{i}\}_{i=1}^{N}$, and obtain their corresponding rewards $\{r_{s}^{i}\}_{i=1}^{N}$, where $r_{s}^{i}=D(y_{s}^{i})$. The advantage of the $i$-th response can be calculated with:
@@ A.3 Automatic Evaluation Details
The sampling temperature is set to $0.8$ and model response length is set to 1536 tokens, same as in training. We use the prompt wrapper in Figure 7 to construct prompts. We use the prompt in Figure 8 for GPT-4o feedback following [minillm].

### v3

@@ A.1 Implement GAD with GRPO
We implement policy optimization of the student with GRPO [27]. We use $q_{G}$ to denote the output distribution of student $G$. For each input prompt $x$, we sample a group of $N$ student responses $\{y_{s}^{i}\}_{i=1}^{N}$, and obtain their corresponding rewards $\{r_{s}^{i}\}_{i=1}^{N}$, where $r_{s}^{i}=D(y_{s}^{i})$. The advantage of the $i$-th response can be calculated with:
@@ A.3 Automatic Evaluation Details
We use greedy sampling and the model response length is set to 1536 tokens. We use the prompt wrapper in Figure 7 to construct prompts. We use the prompt in Figure 8 for GPT-4o feedback following [10]. The reported GPT-4o score is defined as the student’s score divided by the sum of the student’s score and the reference answer’s score.